# Learning hidden cascades via classification

### v2

@@ 4.2 Evaluation on Empirical Social Graph (Insiders Network)
[…]
@@ 5.1 Insiders’ Data
Individuals who trade in a specific stock are referred to as agents, and the network formed through their social connections is the insiders network. This network includes agents who serve as board members of one or more companies. These individuals are believed to possess valuable insider information regarding the future direction of stock prices goel2024 . Those who trade based on such information reportedly achieve substantial returns—approximately 35% over a 21-day period Ahern2017 . In our empirical analysis, we assume that information disseminates within this network prior to the release of public company announcements, triggering a cascade of opportunistic trading. When insider information becomes available, agents are assumed to act on it by executing profitable buy or sell trades. The hypothesis is that information may begin to spread from a company’s board through social links prior to a public announcement. To analyze this hypothesis, we utilize a unique and comprehensive dataset compiled from multiple sources. i) Insider network data in Finland: Being an insider in the same company establishes a social link among all members. Overlapping board memberships consequently form a large, interconnected social network. ii) Board member information and insider trading disclosures: Mandatory disclosure notifications by insiders enable us to identify partial trading patterns. iii) Pseudonymized trading data: A unique and extensive dataset containing detailed trading records of all investors across all securities listed on the Helsinki Stock Exchange. iv) Company announcement data: Information regarding the timing and nature of public company disclosures (See Tables 5 and 6 in Appendix  E for descriptive statistics).
@@ 5.2 Results
In the context of the insider trading network, each spreading cascade corresponds to an information cascade that precedes a public announcement. For a company issuing $n$ public announcements, this results in $n$ independent information cascades. Within each cascade, the participating entities are investors, and their observable symptoms are defined based on trading outcomes: a positive symptom indicates a profitable trade, a negative symptom denotes a loss, and the absence of a symptom represents no trading activity. These outcomes collectively form a trade matrix, which captures investor behavior across cascades in a structured format. The goal is to infer the spreading model’s parameters: $p$ and $q$. Prior to inference, it is necessary to extract the ground truth feature vectors from the trade data. Algorithm 4 in Appendix B details the procedure used to construct the feature set $\mathcal{F}_{\text{GT}}$. All subsequent experiments adhere to the setup described in Appendix A. For each company, the $p$ and $q$ probabilities are estimated separately for the announcement and non-announcement periods.
Table 8 (see Appendix E) presents the average inferred parameters ($\hat{p}$ and $\hat{q}$) across a set of companies, comparing the announcement and non-announcement periods. The non-announcement window is intended to serve as a baseline, under the assumption that information transmission within insider networks should be minimal in the absence of public news. If investors made their decisions completely independently, the parameters $p$ and $q$ during these periods should be close to zero. However, the data reveals a different story. Both $\hat{p}_{n}$ and $\hat{q}_{n}$ are frequently non-zero and, in several cases, considerably high. This is because the baseline model does not account for investors’ collective responses to exogenous factors, such as macroeconomic news and market volatility. These factors can largely explain the observed co-occurrences in their trading activity and trigger synchronized trades across the stock market baltakiene2021identification . As a consequence, the parameter estimates $\hat{p}$ are expected to be positive even outside of the announcement periods, serving as a useful comparative reference for evaluating parameter estimates between announcement and non-announcement periods.
@@ 6 Conclusion
Limitations and Discussion
This study introduces a general framework for estimating the spreading model’s parameters in networks based on partially observed cascades. Although we apply the method to insider trading data, the underlying model is broadly applicable beyond financial contexts. A key avenue for future work is inferring node-specific probabilities ($\hat{p}$, $\hat{q}$) to capture heterogeneity in how individual entities transmit or respond to information. In the Insider trading context, this would allow for more precise identification of individuals involved in early information access or dissemination, helping streamline regulatory investigations and reduce monitoring efforts. One limitation is that we evaluate the framework solely under the Independent Cascade (IC) model. While the IC model is interpretable and widely used, validating the approach under alternative diffusion processes, such as Linear Threshold or continuous-time models, remains an open direction. Further, identifying the exact source of a cascade is not addressed in this work and remains a potential direction for future research.
Acknowledgements
The First author and the corresponding author acknowledge support from the Foundation for the Advancement of Finnish Securities Markets under the project “Tip Chains: Modeling and Analyzing Information Flows in Insider Networks” (Grant Agreement No. 20220021). The corresponding author acknowledges support from the European Union’s Horizon Europe programme under the Marie Skłodowska-Curie Actions (Grant Agreement No. 101150609, Project: HiddenTipChains).
@@ Appendix A Experiment Setup
      Mark all the edges (transmission links) as "not yet tried." This means that these edges have not been used for propagation.
@@
        If the edge state is "not yet tried," a random number $r$ is generated in the range $[0,1]$. The neighboring entity is set as infected if the edge weight $P$ is greater than $r$, and the edge state is then updated to "tried." This indicates that the edge has been used for propagation and cannot be utilized again.
@@
Appendix B Algorithms

### v3

@@ 4.2 Evaluation on Empirical Social Graph (Insiders Network)
The proposed method is subsequently validated using the insiders’ network, which is described in detail in Section 5.1. In this setup, rather than generating ground truth parameters $\bm{\theta}$ from transaction data, we fix the ground truth parameters a priori. The objective is to evaluate the model’s effectiveness in a larger, more complex network with a greater number of entities. Unlike previous setups, this experiment involves multiple seed entities. The spreading process begins from each seed entity sequentially, with an incremental time offset. Specifically, the first seed entity initiates spreading at time $t$, the second at time $t+1$, and so on. As shown in Figure 3(c) in Appendix F, the nodes marked in red represent these seed entities, which are distributed across different regions of the network to ensure extensive spread.
4.3 Benchmarking, Robustness and Efficiency analysis
First, we benchmarked our method against four alternative approaches in Appendix B: i) the approach proposed by Gutmann et al. (2018) and three Graph Neural Network Architectures. Our distribution-matching approach consistently achieved lower errors and more stable parameter recovery. To evaluate the framework’s performance and efficiency, we conducted several robustness checks in Appendix C. First, we relax the assumption of a uniform propagation probability $p$ and examine robustness under node-specific propagation probabilities. The results, reported in Appendix C.1, show that our method remains effective. Second, in Appendix C.2, we evaluate performance under both the Independent Cascade and Threshold models, and again our approach proves effective. Third, in Appendix C.3 we tested robustness across diverse network topologies, including tree, loopy, and empirical insider trading networks, while varying the spreading model parameters $p$ and $q$ over the set $\{0.1,0.3,0.5,0.7,0.9\}$, learning the corresponding estimates $\hat{p}$ and $\hat{q}$ for each combination, and observed consistent accuracy and parameter recovery across all settings. Finally, to analyze computational efficiency, in Appendix C.4 we evaluated scalability on synthetic BA graphs, and find that runtime grows sublinearly to slightly superlinearly with network size. These results demonstrate that our framework generalizes well, scales efficiently, and remains robust across different settings, diffusion models and network topologies.
@@ 5.1 Insiders’ Data
Individuals who trade in a specific stock are referred to as agents, and the network formed through their social connections is the insiders network. This network includes agents who serve as board members of one or more companies. These individuals are believed to possess valuable insider information regarding the future direction of stock prices Goel et al. (2024). Those who trade based on such information reportedly achieve substantial returns—approximately 35% over a 21-day period Ahern (2017). In our empirical analysis, we assume that information disseminates within this network prior to the release of public company announcements, triggering a cascade of opportunistic trading. When insider information becomes available, agents are assumed to act on it by executing profitable buy or sell trades. The hypothesis is that information may begin to spread from a company’s board through social links prior to a public announcement. To analyze this hypothesis, we utilize a unique and comprehensive dataset compiled from multiple sources. i) Insider network data in Finland: Being an insider in the same company establishes a social link among all members. Overlapping board memberships consequently form a large, interconnected social network. ii) Board member information and insider trading disclosures: Mandatory disclosure notifications by insiders enable us to identify partial trading patterns. iii) Pseudonymized trading data: A unique and extensive dataset containing detailed trading records of all investors across all securities listed on the Helsinki Stock Exchange. iv) Company announcement data: Information regarding the timing and nature of public company disclosures (See Tables 11 and 12 in Appendix  G for descriptive statistics).
@@ 5.2 Results
In the context of the insider trading network, each spreading cascade corresponds to an information cascade that precedes a public announcement. For a company issuing $n$ public announcements, this results in $n$ independent information cascades. Within each cascade, the participating entities are investors, and their observable symptoms are defined based on trading outcomes: a positive symptom indicates a profitable trade, a negative symptom denotes a loss, and the absence of a symptom represents no trading activity. These outcomes collectively form a trade matrix, which captures investor behavior across cascades in a structured format. The goal is to infer the spreading model’s parameters: $p$ and $q$. Prior to inference, it is necessary to extract the ground truth feature vectors from the trade data. Algorithm 4 in Appendix D details the procedure used to construct the feature set $\mathcal{F}_{\text{GT}}$. All subsequent experiments adhere to the setup described in Appendix A. For each company, the $p$ and $q$ probabilities are estimated separately for the announcement and non-announcement periods.
Table 14 (see Appendix G) presents the average inferred parameters ($\hat{p}$ and $\hat{q}$) across a set of companies, comparing the announcement and non-announcement periods. The non-announcement window is intended to serve as a baseline, under the assumption that information transmission within insider networks should be minimal in the absence of public news. If investors made their decisions completely independently, the parameters $p$ and $q$ during these periods should be close to zero. However, the data reveals a different story. Both $\hat{p}_{n}$ and $\hat{q}_{n}$ are frequently non-zero and, in several cases, considerably high. This is because the baseline model does not account for investors’ collective responses to exogenous factors, such as macroeconomic news and market volatility. These factors can largely explain the observed co-occurrences in their trading activity and trigger synchronized trades across the stock market Baltakiene et al. (2021). As a consequence, the parameter estimates $\hat{p}$ are expected to be positive even outside of the announcement periods, serving as a useful comparative reference for evaluating parameter estimates between announcement and non-announcement periods.
@@ Appendix A Experiment Setup
      Mark all the edges (transmission links) as ”not yet tried.” This means that these edges have not been used for propagation.
@@
        If the edge state is ”not yet tried,” a random number $r$ is generated in the range $[0,1]$. The neighboring entity is set as infected if the edge weight $P$ is greater than $r$, and the edge state is then updated to ”tried.” This indicates that the edge has been used for propagation and cannot be utilized again.
@@
Appendix B Benchmarking Against Alternative Models
We benchmark our framework against the method proposed in Gutmann et al. (2018), which uses Approximate Bayesian Computation (ABC) for likelihood-free inference. A key feature of this approach is that the discrepancy function is defined independently of the internal cascade dynamics, making it applicable in settings with limited temporal information. To ensure a fair comparison, we implemented Gutmann et al. (2018) under the same experimental conditions as our framework. The results in Table 2 demonstrate that Gutmann et al. (2018) performs poorly in this setting, with the estimated propagation probabilities ($\hat{p}$) and symptom probabilities ($\hat{q}$) exhibiting substantial variance and notable deviations from the ground-truth values. The issue is particularly pronounced in Loopy graphs, where the Mean Squared Error (MSE) reaches $2.56\times 10^{-1}$, underscoring the method’s sensitivity to noisy observations. In contrast, our framework consistently yields parameter estimates close to the true values and achieves markedly lower MSE across both Tree and Loopy structures (see Table 7 and 8 ).
Table 2: Performance of the method in Gutmann et al. (2018) under the IC model with symptom probability $q=0.7$. Results are shown for Tree and Loopy network topologies. Here, $p$ denotes the true propagation probability, and $\hat{p}$, $\hat{q}$ are the estimated parameters.
Tree Loopy
$p$ $\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$ $\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 $0.33\pm 0.141$ $0.46\pm 0.299$ $9.11\times 10^{-2}$ 0.35 $0.28\pm 0.128$ $0.58\pm 0.214$ $5.23\times 10^{-2}$ 0.37
0.3 $0.25\pm 0.027$ $0.43\pm 0.224$ $5.55\times 10^{-2}$ 0.38 $0.29\pm 0.092$ $0.53\pm 0.221$ $4.18\times 10^{-2}$ 0.38
0.5 $0.73\pm 0.129$ $0.69\pm 0.006$ $3.27\times 10^{-2}$ 0.40 $0.71\pm 0.107$ $0.69\pm 0.022$ $2.65\times 10^{-2}$ 0.42
0.7 $0.63\pm 0.097$ $0.70\pm 0.006$ $5.38\times 10^{-3}$ 0.36 $0.42\pm 0.365$ $0.51\pm 0.345$ $1.42\times 10^{-1}$ 0.59
0.9 $0.62\pm 0.230$ $0.63\pm 0.197$ $8.57\times 10^{-2}$ 0.46 $0.41\pm 0.356$ $0.45\pm 0.318$ $2.56\times 10^{-1}$ 0.64
We additionally benchmark our model with several well-known Graph Neural Network (GNN) variants, including Graph Attention Network(GAT), alongside Graph Convolutional Network(GCN)-based architectures, and report the best results. While GNN-based approaches offer considerable modeling flexibility, their effective deployment involves non-trivial design choices such as selecting the appropriate GNN variant, tuning embedding dimensions, and determining network depth. These steps are computationally demanding and require substantial amounts of data for training.
For comparison, we implemented a two-layer GAT, two-layer GCN and four-layer GCN and evaluated it under the same experimental setting. The model is trained over the grid $[0,1]$ with a step size of 0.1 for both $p$ and $q$, and its architecture can be expressed as:
The GCN and GAT architectures used in our experiments are summarized in the Table 3. Each model maps node features $x$ to the estimated propagation and symptom probabilities $\hat{p},\hat{q}$].
Table 3: Comparison of Graph Neural Network Architectures.
Architecture Layers Activation Pooling Output
2-layer GAT GAT $\rightarrow$ GAT ReLU GlobalMeanPool Sigmoid
2-layer GCN GCN $\rightarrow$ GCN ReLU GlobalMeanPool Sigmoid
4-layer GCN GCN $\rightarrow$ GCN $\rightarrow$ GCN $\rightarrow$ GCN ReLU GlobalMeanPool Sigmoid
Across all variants, the GNN-based models underperformed relative to our distribution-matching (DC) approach, particularly for Loopy graphs, where high standard deviation and elevated mean squared error (MSE) were observed. In contrast, our framework delivers robust parameter estimates without requiring supervised training, making it significantly more practical for data-scarce scenarios. Detailed numerical results are reported in Table 4.
These observations justify our methodological choice: while GNNs present an interesting alternative, the additional complexity and training requirements did not yield improved performance in our experimental setting. Our distribution-matching framework aligns the feature distributions of observed and simulated data, eliminating the need for large labeled datasets while maintaining reliable performance. Exploring advanced GNN architectures and training strategies remains a promising avenue for future research.
Table 4: Comparison of parameter estimation across multiple models (GAT, GCN, and GCN 4-layer) for Tree and Loopy graphs under the IC model ($q=0.7$ fixed). Each cell shows mean ± standard deviation over 5 samples, and MSE is reported in scientific notation.
Tree Loopy
$p$ $\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\hat{p}$ $\hat{q}$ $\overline{MSE}$
2-layer GAT
0.1 $0.20\pm 0.020$ $0.67\pm 0.035$ $6.69\times 10^{-3}$ $0.13\pm 0.006$ $0.44\pm 0.087$ $3.81\times 10^{-2}$
0.3 $0.28\pm 0.042$ $0.73\pm 0.039$ $2.40\times 10^{-3}$ $0.32\pm 0.011$ $0.62\pm 0.025$ $3.58\times 10^{-3}$
0.5 $0.47\pm 0.042$ $0.81\pm 0.025$ $8.22\times 10^{-3}$ $0.51\pm 0.028$ $0.65\pm 0.004$ $1.54\times 10^{-3}$
0.7 $0.72\pm 0.012$ $0.84\pm 0.070$ $1.29\times 10^{-2}$ $0.68\pm 0.023$ $0.67\pm 0.004$ $1.03\times 10^{-3}$
0.9 $0.91\pm 0.008$ $0.73\pm 0.005$ $4.16\times 10^{-4}$ $0.89\pm 0.006$ $0.65\pm 0.010$ $1.14\times 10^{-3}$
2-layer GCN
0.1 $0.17\pm 0.034$ $0.59\pm 0.094$ $1.36\times 10^{-2}$ $0.08\pm 0.006$ $0.67\pm 0.040$ $1.50\times 10^{-3}$
0.3 $0.33\pm 0.055$ $0.70\pm 0.057$ $3.56\times 10^{-3}$ $0.29\pm 0.007$ $0.70\pm 0.014$ $1.56\times 10^{-4}$
0.5 $0.50\pm 0.048$ $0.73\pm 0.047$ $2.78\times 10^{-3}$ $0.49\pm 0.007$ $0.70\pm 0.009$ $1.09\times 10^{-4}$
0.7 $0.71\pm 0.012$ $0.72\pm 0.034$ $9.16\times 10^{-4}$ $0.66\pm 0.011$ $0.70\pm 0.011$ $1.04\times 10^{-3}$
0.9 $0.91\pm 0.009$ $0.68\pm 0.013$ $4.96\times 10^{-4}$ $0.90\pm 0.005$ $0.69\pm 0.004$ $9.50\times 10^{-5}$
4-layer GCN
0.1 $0.17\pm 0.045$ $0.41\pm 0.104$ $4.98\times 10^{-2}$ $0.06\pm 0.002$ $0.75\pm 0.063$ $3.93\times 10^{-3}$
0.3 $0.23\pm 0.016$ $0.82\pm 0.025$ $9.81\times 10^{-3}$ $0.29\pm 0.013$ $0.66\pm 0.011$ $1.20\times 10^{-3}$
0.5 $0.45\pm 0.046$ $0.75\pm 0.062$ $5.62\times 10^{-3}$ $0.45\pm 0.034$ $0.69\pm 0.011$ $1.96\times 10^{-3}$
0.7 $0.67\pm 0.007$ $0.72\pm 0.027$ $9.96\times 10^{-4}$ $0.71\pm 0.059$ $0.69\pm 0.006$ $1.82\times 10^{-3}$
0.9 $0.89\pm 0.012$ $0.74\pm 0.006$ $1.07\times 10^{-3}$ $0.91\pm 0.004$ $0.69\pm 0.008$ $1.47\times 10^{-4}$
Appendix C Robustness and Efficiency Analysis
C.1 Node-Specific Propagation Probabilities
We conducted robustness experiments by sampling $p\sim U(0,1)$ with fixed $q$ and further compared our framework against a two-layer GCN model that outputs node-specific $p$. Table 5 summarizes the results across three propagation probability ranges: $(0,1)$, $(0,0.5)$, and $(0.5,1)$, with actual $p$ uniformly distributed in each range. The GCN performs reasonably well for $p\in(0,0.5)$ but its accuracy deteriorates sharply for higher propagation values, failing almost entirely when $p\in(0.5,1)$. In contrast, DC maintains stable accuracy and low error across all ranges, demonstrating robustness under both low and high propagation probabilities. The slightly lower accuracy and higher MSE observed for DC in the $(0,1)$ case are expected, as this setting spans the full propagation range and is inherently more challenging than the narrower intervals. The number of assignments was kept consistent across all settings. These findings reveal a key limitation of GCN-based models in handling heterogeneous propagation dynamics, whereas the proposed distribution-matching approach remains effective across regimes. Nevertheless, due to the stochastic nature of diffusion processes, perfectly estimating node-specific $p$ or $q$ values is not feasible.
Table 5: Performance of DC vs. GCN models ($q=0.7$) under different propagation probability ranges. Metrics: MSE, Acc@0.1, Acc@0.2. Acc@0.1 measures the proportion of nodes whose predicted parameter lies within $\pm 0.1$ of the true value (similarly for Acc@0.2). Each GCN experiment uses 1000 assignments with an 80-10-10 split.
Range of $p$ and $\hat{p}$ Model MSE Acc@0.1 Acc@0.2
$(0,1)$ GCN 0.1444 0.20 0.39
DC 0.1401 0.25 0.47
$(0,0.5)$ GCN 0.0328 0.39 0.69
DC 0.0362 0.41 0.69
$(0.5,1)$ GCN 0.5274 0.01 0.06
DC 0.0362 0.41 0.69
C.2 Robustness to Diffusion Model Variations
Although our main experiments focused on the IC model, the proposed learning framework is model-agnostic and can be applied with any diffusion process, provided the propagation dynamics are specified. To illustrate this, we additionally evaluated our method under the Linear Threshold model Granovetter (1978). The results under LT closely mirror those obtained for IC, yielding comparable parameter recovery and predictive performance (see Table 6). These findings confirm that our approach remains robust across different diffusion mechanisms, while exploration of further models is left for future work.
Table 6: Evaluation under the LT model with symptom probability $q=0.7$. Here, $p_{thr}$ denotes the activation threshold (fraction of neighbors required for activation), and $\hat{p}_{thr}$ is its estimated value.
Tree Loopy
$p_{thr}$ $\hat{p}_{thr}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$ $\hat{p}_{thr}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 $0.18\pm 0.007$ $0.78\pm 0.015$ $1.38\times 10^{-2}$ 0.43 $0.19\pm 0.064$ $0.69\pm 0.057$ $3.66\times 10^{-3}$ 0.43
0.3 $0.27\pm 0.035$ $0.61\pm 0.010$ $9.99\times 10^{-3}$ 0.43 $0.30\pm 0.127$ $0.71\pm 0.078$ $1.25\times 10^{-5}$ 0.43
0.5 $0.57\pm 0.028$ $0.70\pm 0.002$ $6.32\times 10^{-3}$ 0.42 $0.60\pm 0.064$ $0.73\pm 0.037$ $4.85\times 10^{-3}$ 0.42
0.7 $0.68\pm 0.026$ $0.70\pm 0.004$ $1.44\times 10^{-3}$ 0.43 $0.79\pm 0.057$ $0.70\pm 0.000$ $4.05\times 10^{-3}$ 0.42
0.9 $0.88\pm 0.039$ $0.70\pm 0.002$ $1.15\times 10^{-3}$ 0.43 $0.86\pm 0.007$ $0.74\pm 0.000$ $1.81\times 10^{-3}$ 0.42
C.3 Robustness to Network Topology
To further examine the stability of our approach, we conducted simulation experiments on synthetic and empirical networks with known ground truth (Tables 7, 8, and 9). These experiments systematically varied the propagation probability $p$ and the symptom probability $q$, and were carried out on three distinct graph structures: tree, loopy, and empirical topologies. Importantly, these evaluations focus exclusively on the structural properties of the networks, independent of investor behavior or transaction data.
Table 7: Robustness Check for Tree Across Multiple Values of $p$ and $q$
$q$ (0.1)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.13 ± 0.040 0.10 ± 0.000 $1.20\times 10^{-3}$ 0.43
0.3 0.29 ± 0.020 0.12 ± 0.020 $7.03\times 10^{-4}$ 0.43
pp 0.5 0.46 ± 0.060 0.10 ± 0.000 $2.22\times 10^{-3}$ 0.43
0.7 0.70 ± 0.005 0.10 ± 0.000 $1.62\times 10^{-5}$ 0.43
0.9 0.90 ± 0.004 0.10 ± 0.009 $6.44\times 10^{-5}$ 0.43
$q$ (0.3)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.15 ± 0.030 0.36 ± 0.140 $1.17\times 10^{-2}$ 0.43
0.28 ± 0.030 0.31 ± 0.020 $7.62\times 10^{-4}$ 0.43
0.49 ± 0.007 0.29 ± 0.020 $2.94\times 10^{-4}$ 0.43
0.70 ± 0.004 0.30 ± 0.005 $3.64\times 10^{-5}$ 0.43
0.90 ± 0.000 0.30 ± 0.000 $5.14\times 10^{-6}$ 0.43
$q$ (0.5)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.12 ± 0.023 0.54 ± 0.069 $3.11\times 10^{-3}$ 0.43
0.3 0.29 ± 0.030 0.51 ± 0.036 $1.05\times 10^{-3}$ 0.43
pp 0.5 0.44 ± 0.044 0.57 ± 0.112 $9.64\times 10^{-3}$ 0.44
0.7 0.68 ± 0.040 0.55 ± 0.112 $6.67\times 10^{-3}$ 0.44
0.9 0.90 ± 0.006 0.50 ± 0.000 $7.73\times 10^{-6}$ 0.43
$q$ (0.7)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.14 ± 0.058 0.68 ± 0.051 $3.37\times 10^{-3}$ 0.43
0.29 ± 0.055 0.74 ± 0.096 $5.58\times 10^{-3}$ 0.43
0.49 ± 0.016 0.69 ± 0.024 $3.70\times 10^{-4}$ 0.43
0.69 ± 0.019 0.73 ± 0.055 $1.76\times 10^{-3}$ 0.43
0.88 ± 0.030 0.74 ± 0.077 $3.10\times 10^{-3}$ 0.47
$q$ (0.9)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.12 ± 0.040 0.73 ± 0.080 $1.71\times 10^{-2}$ 0.43
0.3 0.30 ± 0.010 0.90 ± 0.010 $9.44\times 10^{-5}$ 0.43
pp 0.5 0.51 ± 0.020 0.90 ± 0.070 $2.08\times 10^{-3}$ 0.43
0.7 0.68 ± 0.030 0.92 ± 0.040 $1.18\times 10^{-3}$ 0.49
0.9 0.95 ± 0.050 0.84 ± 0.080 $7.45\times 10^{-3}$ 0.74
$q$ (0.1)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.10 ± 0.005 0.10 ± 0.004 $1.98\times 10^{-5}$ 0.43
pp 0.3 0.30 ± 0.000 0.10 ± 0.005 $1.24\times 10^{-5}$ 0.43
0.5 0.50 ± 0.004 0.10 ± 0.000 $1.23\times 10^{-5}$ 0.44
0.7 0.70 ± 0.005 0.10 ± 0.000 $2.05\times 10^{-5}$ 0.43
0.9 0.90 ± 0.005 0.10 ± 0.000 $3.06\times 10^{-5}$ 0.43
$q$ (0.3)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.11 ± 0.017 0.28 ± 0.019 $4.30\times 10^{-4}$ 0.43
0.30 ± 0.009 0.31 ± 0.005 $4.79\times 10^{-5}$ 0.43
0.50 ± 0.006 0.30 ± 0.000 $1.40\times 10^{-5}$ 0.43
0.70 ± 0.000 0.30 ± 0.000 $3.45\times 10^{-6}$ 0.43
0.89 ± 0.014 0.30 ± 0.000 $6.73\times 10^{-5}$ 0.43
$q$ (0.5)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.09 ± 0.015 0.53 ± 0.052 $1.78\times 10^{-3}$ 0.43
0.3 0.30 ± 0.005 0.50 ± 0.017 $1.44\times 10^{-4}$ 0.48
pp 0.5 0.50 ± 0.006 0.50 ± 0.000 $9.17\times 10^{-6}$ 0.43
0.7 0.70 ± 0.000 0.50 ± 0.000 $6.93\times 10^{-6}$ 0.43
0.9 0.90 ± 0.005 0.50 ± 0.000 $1.33\times 10^{-5}$ 0.43
$q$ (0.7)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.11 ± 0.023 0.70 ± 0.057 $1.69\times 10^{-3}$ 0.43
0.30 ± 0.000 0.70 ± 0.010 $7.91\times 10^{-5}$ 0.42
0.50 ± 0.000 0.70 ± 0.000 $4.06\times 10^{-6}$ 0.43
0.70 ± 0.000 0.70 ± 0.000 $2.37\times 10^{-6}$ 0.43
0.90 ± 0.007 0.70 ± 0.000 $2.27\times 10^{-5}$ 0.43
$q$ (0.9)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.11 ± 0.021 0.84 ± 0.109 $6.74\times 10^{-3}$ 0.44
0.3 0.31 ± 0.005 0.87 ± 0.025 $9.10\times 10^{-4}$ 0.43
pp 0.5 0.50 ± 0.000 0.90 ± 0.000 $2.76\times 10^{-7}$ 0.43
0.7 0.70 ± 0.000 0.90 ± 0.000 $1.35\times 10^{-6}$ 0.43
0.9 0.91 ± 0.008 0.90 ± 0.000 $4.50\times 10^{-5}$ 0.43
$q$ (0.1)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.10 ± 0.009 0.10 ± 0.005 $7.43\times 10^{-5}$ $1.62\times 10^{-2}$
0.3 0.37 ± 0.027 0.10 ± 0.009 $2.66\times 10^{-3}$ $1.07\times 10^{-2}$
pp 0.5 0.54 ± 0.024 0.10 ± 0.005 $8.67\times 10^{-4}$ $6.04\times 10^{-3}$
0.7 0.71 ± 0.007 0.10 ± 0.008 $1.07\times 10^{-4}$ $4.83\times 10^{-3}$
0.9 0.89 ± 0.022 0.10 ± 0.005 $2.21\times 10^{-4}$ $4.75\times 10^{-3}$
$q$ (0.3)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.11 ± 0.012 0.30 ± 0.009 $1.35\times 10^{-4}$ $1.66\times 10^{-2}$
0.38 ± 0.041 0.30 ± 0.010 $3.74\times 10^{-3}$ $1.16\times 10^{-2}$
0.55 ± 0.059 0.30 ± 0.007 $2.69\times 10^{-3}$ $6.96\times 10^{-3}$
0.72 ± 0.027 0.30 ± 0.008 $5.98\times 10^{-4}$ $6.45\times 10^{-3}$
0.91 ± 0.025 0.30 ± 0.011 $3.45\times 10^{-4}$ $5.88\times 10^{-3}$
$q$ (0.5)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.11 ± 0.013 0.49 ± 0.008 $1.16\times 10^{-4}$ $1.68\times 10^{-2}$
0.3 0.36 ± 0.033 0.50 ± 0.004 $2.43\times 10^{-3}$ $1.18\times 10^{-2}$
pp 0.5 0.56 ± 0.036 0.50 ± 0.005 $2.36\times 10^{-3}$ $7.89\times 10^{-3}$
0.7 0.75 ± 0.019 0.50 ± 0.011 $1.34\times 10^{-3}$ $7.08\times 10^{-3}$
0.9 0.91 ± 0.036 0.50 ± 0.000 $6.06\times 10^{-4}$ $6.57\times 10^{-3}$
$q$ (0.7)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.11 ± 0.004 0.69 ± 0.016 $2.80\times 10^{-4}$ $1.67\times 10^{-2}$
0.37 ± 0.019 0.69 ± 0.005 $2.53\times 10^{-3}$ $1.18\times 10^{-2}$
0.53 ± 0.021 0.70 ± 0.007 $7.28\times 10^{-4}$ $7.12\times 10^{-3}$
0.72 ± 0.027 0.70 ± 0.015 $5.41\times 10^{-4}$ $6.71\times 10^{-3}$
0.91 ± 0.046 0.70 ± 0.005 $8.15\times 10^{-4}$ $6.34\times 10^{-3}$
$q$ (0.9)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.11 ± 0.005 0.90 ± 0.011 $7.49\times 10^{-5}$ $1.65\times 10^{-2}$
0.3 0.33 ± 0.013 0.90 ± 0.008 $5.44\times 10^{-4}$ $1.06\times 10^{-2}$
p 0.5 0.51 ± 0.004 0.90 ± 0.005 $6.42\times 10^{-5}$ $5.89\times 10^{-3}$
0.7 0.72 ± 0.019 0.91 ± 0.005 $5.01\times 10^{-4}$ $5.44\times 10^{-3}$
0.9 0.91 ± 0.022 0.91 ± 0.005 $2.74\times 10^{-4}$ $4.70\times 10^{-3}$
C.4 Computational Efficiency and Scalability
The proposed framework is highly parallelizable, as agent-level classification tasks are independent, allowing efficient utilization of computational resources. Scalability was evaluated through synthetic experiments on Barabási–Albert (BA) graphs of varying sizes (up to 10,000 nodes, parameter $m=2$). As shown in Table 10, DC exhibits sublinear to slightly superlinear growth, with empirical complexity approximately $T(n)\approx k\times n^{0.5-0.65}$. For a 10,000-node BA graph, the average runtime was approximately 1,279 s under parallel execution. By comparison, the GCN baseline required over 6,200 s for a 1,000-node graph and is projected to exceed 122,000 s for 10,000 nodes, making it impractical for large-scale scenarios. These results highlight that DC scales efficiently to large networks while substantially reducing computational cost compared to deep learning–based alternatives.
Table 10: Comparison of execution time (in seconds) between our method DC and the GCN baseline on tree-structured networks, using a MacBook M1 Pro with parallel execution on 10 CPU cores.
Network Size DC (Avg $\pm$ Std) GCN (Avg $\pm$ Std)
100 $89.9\pm 3.2$ $169.2\pm 2.5$
1000 $289.6\pm 11.8$ $6299.9\pm 275.0$
10000 $1279.0\pm 47.7$ $\sim 122{,}000$ (estimated)
Appendix D Algorithms